1. Data set basics Elements, variables, observations, measurements, and data types An analyst collected data on some of the specifications for new vehicles produced for the 2004 model year The top and bottom portions of the data set are reproduced below; dots indicate that the intervening rows in the data Aa Aa set are not displayed. (Source: Roger W. Johnson, Journal of Statistics Education Data Archive) Vehicle ID Drive Train Price $11,690 $12,585 $14,610 $14,810 $16,385 Engine Size (Liters) 1.6 2.2 2.2 2.2 City Miles per Gallon 28 28 26 26 26 Weight (Pounds) 2,370 2,348 2,617 2,676 2,617 4 424 425 426 427 428 5.6 $26,650 $24,520 $12,800 $16,495 $25,935 5,287 3,485 2,750 3,925 4,435 14 0 2.5 21 16 14 Note that the variable Drive Train has three different values, which are coded as 0 = all-wheel drive, 1-front-wheel drive, and 2 rear-wheel drive Complete the table below to indicate the number of elements, variables, observations, and data values that are included in the complete data set Data Set Term Elements Variables Observations Measurements How Many? Which variables in the data set are categorical? O Price, Engine Size, City Miles per Gallon, and Weight None of the variables All of the variables O Drive Train
Answer #1

number of elements =428

variables =5

observations =428 ( 1 observation set for each element)

measurements =428*5 =2140

categorical variables: Drive train